OpenStudy (anonymous):

Solve the system by using elimination or elimination with multiplication. One point for the steps and one point for the correct solution. 3x + 4y = -2 and 2x - 3y = 10

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

well solve for y in one and plug it into y in the other that is how i would do it.

OpenStudy (anonymous):

3x + 4y = -2 and 2x - 3y = 10 y = -(3/4)*x-1/2 2x-3*(-(3/4)*x-1/2)=10

OpenStudy (anonymous):

thus x=2 plug that back into 3x+4y=-2 3*2+4y=-1 sfy y = -7/4 , x=2